OpenStudy (anonymous):

1.Write an equation in point-slope form of the line through point J(4, 1) with slope –4. 2.Write the equation for the horizontal line that contains point G(–8, 8).

OpenStudy (anonymous):

y-1=-4(x-4)

OpenStudy (anonymous):

and the other? please

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

 the answer depends on how u take ur horizontal axis....

OpenStudy (anonymous):

im still confussed sorriee

OpenStudy (anonymous):

my choices are y = –8 x = 8 y = 1x + 8 y = 8